OpenStudy (routohan75):

What is the slope, and what is the y-intercept of the line? y=2/3x-4 slope = y-intercept =

OpenStudy (routohan75):



OpenStudy (anonymous):

y=mx+b m is your slope and b is your y intercept

OpenStudy (anonymous):

Slope: 23 Y-Intercept: −4

OpenStudy (anonymous):

2/3

OpenStudy (andrewishere):

I agree. 2/3. y=mx+b, where m + slope and b = y-intercept. Since we are talking about slope, we are talking about m, which is represented by 2/3, so the slope is 2/3. :)

OpenStudy (routohan75):

Ow ok I think I am starting to get it. FINALLY!!